I hope you can help with this word problem. I don't know where to begin and I cannot find an example in the book to go by. Thank you
Problem
Time to swing
The speed T (time in seconds for one complete cycle) of a pendulum is related to the length L(in feet) of the pendulum by the formula 8T^2 = pi^2L. If a child is on a swing with a 10-foot chain, then how long does it take to complete one cycle of the swing.
Hope someone can help please.

Answer by checkley75(3666)

8T^2=pi^2L WHEN L=10 FEET THE THE TIME IS:
8T^2=3.14^2*10
8T^2=9.8596*10
8T^2=98.596
T^2=98.596/8
T^2=12.3245
T=SQRT12.3245
T=3.51 SECONDS TO COMPLETE ONE CYCLE (SWING)